Prove that if N is a submodule of a finitely generated module M over a ring R .Then M/N is also finitely generated R module.


1
Expert's answer
2021-06-15T12:01:04-0400

Suppose "M" is generated by {a1,a2,...,an}.

Let J = <ai + N | 1 ≤ i ≤ n>.

Clearly, J is finitely generated.

We show that M/N = J.

Suppose "y \in J" . Then y = ak + N where ak "\in" M for some "k \in" {1,2,...,n}.

So "y \in M/N" . Hence "J \subset M/N"

Conversely, suppose "z \in M/N." Then "z=x+N" where "x\in M" .

Since "M" is finitely generated, "\exist" ri"\in R" such that "x=\Sigma" riai.

Consider "z=x+N
\\=(\Sigma r_i a_i ) + N
\\=\Sigma(r_i a_i + N) = \Sigma r_i (a_i + N)."

Thus "J""=M\/N."